OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). 5x + y = 5 3x + 2y = 3

OpenStudy (texaschic101):

5x + y = 5 ---> y = 5 - 5x now sub 5 - 5x in for y in the second equation... 3x + 2y = 3 3x + 2(5 - 5x) = 3 now distribute the 2 through the parenthesis. Then combine like terms. Get the x on one side and everything else on the other side. Then solve for x. Can you take it from here ?